OpenStudy (anonymous):

Can somebody please explain how my solutions manual got their answer, for the following problem? An artist intends to paint a 60 square foot mural, on a large wall that measures 18ft by 11ft. He plans to leave a uniform border, unpainted around it, on the wall. The illustration has 'w' standing for the white border around it. The solutions manual got 12 by 5 ft, for the mural. Can somebody explain how they got this?

OpenStudy (anonymous):

if you go through the numbers that multiply to 60 (6x10, 5x12, 3x20), the only one that leaves you with equal space around the sides is 5x12. it's a logic-based answer. hope that helped.

OpenStudy (anonymous):

Not really...you see, I'm supposed to be coming up with an equation for the thing, then using the zero-factor property to factor it out...I've done the solutions manual equation and used the zero-factor property to factor it out...and I got what they got, from the equation...that being 23/3 or 3. I just don't see how they got 12ft by 5ft from 23/3 and 3...

OpenStudy (anonymous):

F9

OpenStudy (anonymous):

F9?

OpenStudy (anonymous):

FINE ANS NO EXPLANATION BY STEPS!!!!!!!!!!

OpenStudy (anonymous):

? There has to be an explanation...it's in the solutions manual..somehow they got from x =23/3 or x=3, to 12ft by 5ft.

OpenStudy (anonymous):

maybe \[(18-2x)(11-2x)=60\]

OpenStudy (anonymous):

Satellite73, that is the equation they used and I used...

OpenStudy (anonymous):

ok then why does that not give the answer?

OpenStudy (anonymous):

Because when you use the zero-factor property on that equation...

OpenStudy (anonymous):

multiply out and get \[4 x^2-58 x+138 = 0\]

OpenStudy (anonymous):

oooooooooooooooooh

OpenStudy (anonymous):

you don't get 12 by 5 ft...

OpenStudy (anonymous):

i see. you cannot use the "zero property" unless the thing is set to 0!

OpenStudy (anonymous):

which I did...

OpenStudy (anonymous):

you cannot use it in the form \[(18-2x)(11-2x)=60\] you first have to write what i did

OpenStudy (anonymous):

I set the thing to 0, and got this equation...(18 - 2w)(11 - 2w) -60 = 0...

OpenStudy (anonymous):

now factor as \[2(x-3)(2x-23)=0\] and now you can use the "zero property"

OpenStudy (anonymous):

I did factor it out as that..

OpenStudy (anonymous):

that is what they mean by "use the zero property". then mean set = 0, then factor, then get your answer. you have \[4x^2-58x+138=0\] i would divide by 2 to get \[2x^2-28x+69=0\] to make factoring easier.

OpenStudy (anonymous):

ok so your choice of answers are \[x=3\] or \[x=\frac{23}{2}=11.5\] and the second one is certainly not right.

OpenStudy (anonymous):

The solutions manual got x = 23/2 or x =3...then said that the answer to the problem is 12 by 5 ft. It did not explain how it got from 23.2 or 3 to 12 by 5 ft...

OpenStudy (anonymous):

so the boarder is 3 feet yes?

OpenStudy (anonymous):

and if one dimension is 18 and you take away 3 from both sides you are left with 12

OpenStudy (anonymous):

the other dimension is 11 and you take away 3 from both sides you get 5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

that is how they got the answer. the solution 23/2 = 11.5 makes no sense in this case.

OpenStudy (anonymous):

An artist intends to paint a 60 square foot mural, on a large wall that measures 18ft by 11ft. we have found that the boarder is 3 feet yes?

OpenStudy (anonymous):

3 feet all around. so you have a rectangle inside a rectangle. the outer one is 18 by 11 the inner one is 3 feet less on both sides, making it...